Choose the slope-intercept form of the equation of the line that passes through (-2,4) and is perdendicular to the graph of 2x+y=-20
Answer by Alan3354(69443)

Choose the slope-intercept form of the equation of the line that passes through (-2,4) and is perdendicular to the graph of 2x+y=-20
--------------
It's a 2 step process.
1st, find the slope, m, of 2x+y=20. Do that by putting it in the slope-intercept form, which means solve for y.
y = -2x + 20
m = -2
The slope of lines perpendicular will be the negative inverse, or +1/2 (ok, 3 steps?)
--------
Step 2.5
Use y-y1 = m*(x-x1) where (x1,y1) is the point (-2,4)
y-4 = (1/2)*(x+2)
y = (1/2)x + 5 slope-intercept form